# Closed-Form Power and Sample Size Calculations for Bayes Factors

### v1

@@ 4 Sample size determination
Both power functions (3) and (5) are straightforward to implement and can be used to obtain power curves as a function of the sample size, or of other parameters. We provide R implementations of both in our package bfpwr (see Appendix A for an illustration). Iterative root-finding can then be applied to determine the sample size such that compelling evidence is obtained with a desired target power under a specified design prior (as pioneered by Weiss,, 1997). It is worth noting that one can also compute a power curve in favor of the null hypothesis. That is, one can look at one minus the power functions (3) and (5) along with a Bayes factor threshold $k>1$ and fixing the design prior to the null hypothesis ($\mu_{d}=\theta_{0}$, $\tau_{d}=0$). In this way, sample sizes can be determined that guarantee a desired probability of compelling evidence for the null.
We will now investigate situations where the sample size can be obtained in closed-form. As for the distribution of the Bayes factor in the previous section, there is again a distinction between sample size determination for Bayes factors with point analysis priors ($\tau=0$) and normal analysis priors ($\tau>0$), we start again with the former.
@@ 4.1 Bayes factor with point analysis prior
The same formula (9) was also found by Strug et al., (2007) for ‘evidential’ sample size calculations, but is unfortunately not well known. Not surprisingly, the two formulas coincide, since Bayes factors and likelihood ratios – the measure of evidence used in evidential/likelihoodist statistics (see e.g., Edwards,, 1971; Royall,, 1997; Blume,, 2002; Strug,, 2018) – are equivalent when the Bayes factor involves only point hypotheses. Our more general formulas (7) and (8) thus enable ‘hybrid Bayesian/likelihoodist’ design that assumes a likelihoodist analysis but can incorporate prior knowledge and uncertainty via Bayesian design prior, similar to how Bayesian design priors can be used to incorporate parameter uncertainty in the design of frequentist hypothesis tests (see e.g., Grieve,, 2022, for an overview of hybrid Bayesian/frequentist design approaches).
@@
To illustrate formula (9), we now assume that $\hat{\theta}$ is a standardized mean difference with unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}=2$ so that $n$ can be interpreted as the sample size per group (see the third row in Table 2). Table 3 shows the sample size (9) based on an assumed effect size of one ($\Delta_{\mu}=1$). We see, for instance, that a sample size of $n=20$ per group is required to achieve $1-\beta=80\%$ power for a Bayes factor threshold $k=1/10$. If the assumed effect size was smaller, the required sample size would become larger. For example, for a half as large effect size, the sample sizes from Table 3 quadruple, e.g., requiring a sample size of $n=80$ per group to have $1-\beta=80\%$ power for a Bayes factor threshold of $k=1/10$.
@@ 4.2 Bayes factor with normal analysis prior
Finding a sample size formula becomes more difficult when we move from point to normal analysis priors. The technical reason is that when we set the power function (5) equal to a target power and try to solve for the sample size $n$, we have $n$ appearing both in and outside logarithms. This forms a transcendental equation that cannot be solved in terms of elementary functions. We were unable to find a closed-form solution in the general case. However, as we will show now, solutions can, under certain conditions, be expressed in terms of the Lambert W function (Corless et al.,, 1996). The Lambert W function is the function $\mathrm{W}(\cdot)$ that satisfies $\mathrm{W}(x)\exp\{\mathrm{W}(x)\}=x$, and it is therefore sometimes also called ‘product logarithm’. It has many fundamental applications, such as the solution of the Schrödinger equation in quantum-mechanics, and has also previously appeared in the context of Bayes factor hypothesis testing (Pawel and Held,, 2022; Wagenmakers,, 2022; Held et al.,, 2022; Pawel et al.,, 2024).
@@
with $\mathrm{W}_{-1}(\cdot)$ the branch of the Lambert W function that satisfies $\mathrm{W}(x)<-1$ for $y\in(-1/e,0)$ (Corless et al.,, 1996), see Appendix D for details.
@@ 5 Application
5.1 Randomized controlled clinical trial on treatment of influenza
MIST Study Group, (1998) conducted a randomized controlled clinical trial to evaluate the efficacy of the drug ‘zanamivir’ in the treatment of influenza A and B. The primary endpoint of the trial was the time in days to relief of clinically important symptoms of influenza. This outcome was treated as continuous by the investigators, and the parameter of interest was the mean difference in time to relief between the treatment and placebo groups $\theta$. We note that it might be preferable to treat this as a survival outcome and use a (log) hazard ratio parameter to quantify the effect size, but to stay close to the original analysis we will use the continuous outcome approach. The null hypothesis was defined as no difference in efficacy of the treatment ($H_{0}\colon\theta=0$), whereas the alternative hypothesis was defined as a clinically relevant difference of one day ($H_{1}\colon\theta=1$). For sample size calculations, the investigators used a standard deviation of $\sigma=2.75$ estimated from a previous study.
@@ 5.1 Randomized controlled clinical trial on treatment of influenza
While this study was analyzed using frequentist methodology, we will now examine what the power and sample size calculations might look like if the planned analysis were performed using Bayes factors. For doing so, we assume a point analysis prior for the Bayes factor that equals the alternative hypothesis specified by the trial investigators ($\mu=1$ and $\tau=0$), so that the Bayes factor corresponds with a likelihood ratio. Hence, we can use the power function (3) and sample size formula (7) to compute power and sample size under point and normal design priors, which we illustrate in the following.
@@
Figure 3: Power and sample size calculations for MIST trial (MIST Study Group,, 1998). The Bayes factor assumed for the analysis contrasts $H_{0}\colon\theta=0$ to $H_{1}\colon\theta=1$. The power is computed under either a point prior at the assumed effect under the alternative (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange).
@@
Figure 3 shows power curves based on the Bayes factor providing strong evidence in favor of the alternative ($\mathrm{BF}_{01}<1/10$) in the top plot, and based on the Bayes factor providing strong evidence in favor of the null ($\mathrm{BF}_{01}>10$) in the bottom plot. The colors indicate under which design prior the power was computed.
@@
Focusing on the top plot, we can see from the green curve that a sample size of at least $n=217$ per group are required to ensure a target power of $1-\beta=90\%$ assuming that the same point prior is used in the design as for the analysis (i.e., a point prior at $\mu=1$). This number increases to $n=384$ when we move to a design prior that incorporates parameter uncertainty (blue curve), i.e., a prior that is still centered around $\mu=1$ but with a standard deviation of $\tau_{d}=0.25$. Finally, if we look at the orange power curve computed assuming that the null hypothesis is true ($\theta=0$), we can see that the probability of misleading evidence for the alternative when the null hypothesis is actually true is very low and appears to be adequately controlled by conventional standards (i.e., below 5%) for each of the two sample sizes.
@@
Focusing now on the bottom plot, we can see that the sample size $n=217$ also ensures a power of $1-\beta=90\%$ for finding evidence for the null hypothesis. This is due to the symmetric nature of the point versus point hypothesis Bayes factor considered in this example, as swapping the null $\theta_{0}$ and alternative $\mu$ in formula (9) does not change the resulting sample size. Finally, looking at the green and blue curves we can see that the probability of misleading evidence in favor of the null when the data are generated from the non-null design priors seems to be reasonably well controlled (below 5%) across the whole range of sample sizes considered.
@@ 5.2 Comparison to Monte Carlo simulation methods
Schönbrodt and Wagenmakers, (2018) proposed a Monte Carlo simulation approach for power and sample size calculations for Bayes factor analyses, and they provide the R package BFDA (Schönbrodt and Stefan,, 2019) for this purpose. The idea is to simulate data sets under an assumed design prior and sample size, and then analyze each data set with a specified Bayes factor. This results in a distribution of Bayes factors from which the power can be calculated. The simulation is then repeated for other sample sizes until the desired power is achieved. This approach can be used in quite general settings, but can be computationally intensive and comes with Monte Carlo error.
@@
Figure 4: Power and sample size calculation for example adapted from Schönbrodt and Wagenmakers, (2018, p. 133). The Bayes factor assumed for the analysis contrasts $H_{0}\colon\theta=0$ to $H_{1}\colon\theta\neq 0$ with $\theta\mid H_{1}\sim\mathrm{N}(\mu=0,\tau^{2}=1/2)$ prior assigned to the standardized mean difference $\theta$ under the alternative. The power is computed under either a point prior at a medium effect size (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange). For comparison, 1000 Monte Carlo simulations are performed with the BFDA package (Schönbrodt and Stefan,, 2019) to obtain a simulation-based power curve approximation (dashed line with one Monte Carlo standard error band).
@@
We will now look at an adaptation of an example examined in Schönbrodt and Wagenmakers, (2018, p. 133). Suppose we want to test the null hypothesis that a standardized mean difference $\theta$ is zero ($H_{0}\colon\theta=0$) versus the alternative hypothesis that it is different from zero ($H_{1}\colon\theta\neq 0$). We assume a $\theta\mid H_{1}\sim\mathrm{N}(0,1/2)$ analysis prior for the standardized mean difference under the alternative, similar to the Cauchy prior with scale $1/\sqrt{2}$ that was assumed by Schönbrodt and Wagenmakers, (2018). As they did, we also investigate two design priors: either a point prior at $\mu_{d}=0.5$ , which is a convention for a ‘medium‘ standardized mean difference in psychology (Cohen,, 1992), or a normal prior at $\mu_{d}=0.5$ with standard deviation $\tau_{d}=0.1$ to incorporate parameter uncertainty. Figure 4 shows the resulting power curves computed from equation (5), along with the BFDA Monte Carlo simulation approximations (with Monte Carlo standard error bands) for comparison. The closed-form power curves were computed instantly while each simulated power curve took about 10 minutes to compute when parallelized across 10 cores on a modern laptop (Lenovo X1 Gen 8 with Intel i7-1355U CPU and 32GB RAM).
@@
We see that closed-form and simulation curves closely align in all cases, although the former shows Monte Carlo error. Looking at the green curves in the top plot, we see that a sample size of $n=153$ per group is sufficient to achieve a target power of $95\%$ with a Bayes factor threshold of $k=1/6$ under the point design prior. This sample size is slightly larger than the $n=146$ that Schönbrodt and Wagenmakers, (2018) obtained in their calculations, which instead assumed a Cauchy analysis prior. As in the previous example, incorporating parameter uncertainty via a normal design prior increases the required sample size, in this case to $n=211$. Looking at the orange curve in the bottom plot, we see that these sample sizes only have modest power of around $20\%$ and $50\%$, respectively, for obtaining evidence in favor of the true null hypothesis (at a threshold of $k=6$). To obtain a power of $95\%$, a sample size of $n=6691$ per group would be required (not shown in Figure 4). This illustrates the well-known fact that evidence accumulates slower for the null than the alternative when the Bayes factor involves testing a point null against a composite alternative with normal analysis prior (Johnson and Rossell,, 2010). Finally, looking at the orange curve in the top plot and the green/blue curves in the bottom plot, we see that the probability of misleading evidence in favor of the incorrect hypothesis appears to be adequately controlled, as these curves are virtually zero over the entire range of sample sizes.
@@ 6.1 Informed Bayesian t-test
with $t$ the $t$-test statistic, $n$ the effective sample size (the actual number of observations/pairs for one-sample/paired $t$-tests, or $n=1/(1/n_{1}+1/n_{2})$ for two-sample $t$-tests), $\mathrm{T}_{\nu}(\cdot\mid\mu,\tau)$ the location-scale $t$ density with degrees of freedom $\nu$, location $\mu$, and scale $\tau$, and $\mathrm{NCT}_{\nu}(\cdot\mid\lambda)$ the non-central $t$ density with degrees of freedom $\nu$ and non-centrality parameter $\lambda$ (Johnson et al.,, 1995, chapters 28 and 31). The subscript $[a,b]$ denotes truncation of a distribution to the interval from $a$ to $b$, i.e., $f(x)_{[a,b]}=\{f(x)1_{[a,b]}(x)\}/\{F(a)-F(b)\}$. For example, for a one-sided test in positive direction, we have $a=0$ and $b=+\infty$, while for a two-sided test we have $a=-\infty$ and $b=+\infty$. When the prior under the alternative is set to a (scaled) Cauchy distribution ($\kappa=1$, $\mu=0$, $a=-\infty$, $b=+\infty$), the Bayes factor reduces to the ‘Jeffreys-Zellner-Siow’ (JZS) Bayes factor (Jeffreys,, 1961; Zellner and Siow,, 1980), which is often used as a ‘default’ Bayes factor in the social sciences (Rouder et al.,, 2009). Setting other values for these parameters allows data analysts to incorporate directionality, such as a one-sided JZS Bayes factor (Wetzels et al.,, 2009), or prior knowledge about the standardized mean difference, potentially improving the efficiency of the test (Stefan et al.,, 2019). The Bayes factor (12) is also implemented in our package in the function tbf01.
@@
To calculate a sample size that ensures a certain target power, we can again use numerical root-finding. To illustrate this, we now revisit the example from Schönbrodt and Wagenmakers, (2018) which was already re-analyzed in Section 5.2 with a normal analysis prior. We now specify the analysis parameters $\mu=0,\tau=1/\sqrt{2},\kappa=1,a=0,b=+\infty$, i.e., a one-sided JZS Bayes factor with scale $1/\sqrt{2}$ (Wetzels et al.,, 2009), as in the original analysis. Using the previously described iterative approach, we calculate that a sample size of $n=143$ is required under a point design prior at $\mu_{d}=0.5$ to achieve a target power of 95%. This differs only slightly from the simulation-based $n=146$ reported by Schönbrodt and Wagenmakers, (2018), possibly due to Monte Carlo error.

### v2

@@ 4 Sample size determination
Both power functions (3) and (6) are straightforward to implement and can be used to obtain power curves as a function of the sample size, or of other parameters. We provide R implementations of both in our package bfpwr (see Appendix A for an illustration). Iterative root-finding can then be applied to determine the sample size such that compelling evidence is obtained with a desired target power under a specified design prior (as pioneered by Weiss, , 1997). It is important to emphasize again that one can also compute a power curve in favor of the null hypothesis. That is, one can look at one minus the power functions (3) and (6) along with a Bayes factor threshold $k>1$ and fixing the design prior to the null hypothesis ($\mu_{d}=\theta_{0}$, $\tau_{d}=0$). In this way, sample sizes can be determined that ensure a desired probability of compelling evidence for both the null and the alternative.
We will now investigate situations where the sample size can be obtained in closed-form. As for the distribution of the Bayes factor in the previous section, there is again a distinction between sample size determination for Bayes factors with point analysis priors ($\tau=0$) and normal analysis priors ($\tau>0$). We start again with the former.
@@ 4.1 Bayes factor with point analysis prior
The same formula (10) was also found by Strug et al., (2007) for ‘evidential’ sample size calculations, but is unfortunately not well known. Not surprisingly, the two formulas coincide, since Bayes factors and likelihood ratios – the measure of evidence used in evidential/likelihoodist statistics (see e.g., Edwards, , 1971; Royall, , 1997; Blume, , 2002; Strug, , 2018; Cahusac, , 2020) – are equivalent when the Bayes factor involves only point hypotheses. Our more general formulas (8) and (9) thus enable ‘hybrid Bayesian/likelihoodist’ design that assumes a likelihoodist analysis but can incorporate prior knowledge and uncertainty via Bayesian design prior, similar to how Bayesian design priors can be used to incorporate parameter uncertainty in the design of frequentist hypothesis tests (see e.g., Grieve, , 2022, for an overview of hybrid Bayesian/frequentist design approaches).
@@
To illustrate formula (10), we now assume that $\hat{\theta}$ is a standardized mean difference so that the unit variance is $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}=2$ and $n$ can be interpreted as the sample size per group (see the third row in Table 2). Table 3 shows the sample size (10) based on an assumed effect size of one ($\Delta_{\mu}=1$). We see, for instance, that a sample size of $n=20$ per group is required to achieve $1-\beta=80\%$ power for a Bayes factor threshold $k=1/10$. If the assumed effect size was smaller, the required sample size would become larger. For example, for a half as large effect size, the sample sizes from Table 3 quadruple, e.g., requiring a sample size of $n=80$ per group to have $1-\beta=80\%$ power for a Bayes factor threshold of $k=1/10$.
@@ 4.2 Bayes factor with normal analysis prior
Finding a sample size formula becomes more difficult when we move from point to normal analysis priors. The technical reason is that when we set the power function (6) equal to a target power and try to solve for the sample size $n$, we have $n$ appearing both in and outside logarithms. This forms a transcendental equation that cannot be solved in terms of elementary functions. We were unable to find a closed-form solution in the general case. However, as we will show now, solutions can, under certain conditions, be expressed in terms of the Lambert W function (Corless et al., , 1996). The Lambert W function is the function $\mathrm{W}(\cdot)$ that satisfies $\mathrm{W}(x)\exp\{\mathrm{W}(x)\}=x$, and it is therefore sometimes also called ‘product logarithm’. It has many fundamental applications and has also previously appeared in the context of Bayes factor hypothesis testing (Pawel and Held, , 2022; Wagenmakers, , 2022; Held et al., , 2022; Pawel et al., , 2024).
@@
with $\mathrm{W}_{-1}(\cdot)$ the branch of the Lambert W function that satisfies $\mathrm{W}(x)<-1$ for $y\in(-1/e,0)$, see Appendix D for details.
@@ 5 Application
5.1 Randomized controlled clinical trial on mirtazapine in dementia
Banerjee et al., (2021) conducted a randomized controlled clinical trial to assess the effect of the antidepressant mirtazapine on agitated behavior in patients with dementia. The primary outcome of the trial was the Cohen-Mansfield Agitation Inventory (CMAI) score at 12 weeks after treatment start. The CMAI score is commonly used to quantify agitation in dementia and it can take values from 29 to 203. The treatment effect was quantified as the difference in mean CMAI between the mirtazapine and placebo groups $\theta$. The null hypothesis was defined as no mean difference between mirtazapine and placebo ($H_{0}\colon\theta=0$), while the alternative hypothesis was formulated as a mean decrease of six CMAI points ($H_{1}\colon\theta=-6$). With a target power of $1-\beta=80\%$, a two-sided level of $\alpha=0.05$, and an assumed standard deviation of $\sigma=15$ CMAI points, formula (1) gives a required per-group sample size of $n=99$, which the investigators adjusted further upwards to 111 to account for attrition. Contrary to their expectations, however, the trial did not find a statistically significant effect of mirtazapine over placebo (estimated mean difference of $\hat{\theta}=-1.74$ with 95% confidence interval from $-7.17$ to $3.69$ and two-sided $p$-value of $p=0.53$ for the null hypothesis of no difference).
@@ 5.1 Randomized controlled clinical trial on mirtazapine in dementia
While this study was designed and analyzed using frequentist methodology, we will now examine what the analysis, power, and sample size calculations might have looked like if the planned analysis were performed using Bayes factors. To do this, we assume a point analysis prior equal to the alternative hypothesis specified by the trial investigators ($\mu=-6$ and $\tau=0$), so that the Bayes factor corresponds to a likelihood ratio. Based on the estimated mean difference $\hat{\theta}=-1.74$ and its standard error $\sigma_{\hat{\theta}}/\sqrt{n}=2.77$ (recalculated from the confidence interval), we can apply equation (2) and obtain a Bayes factor $\text{BF}_{01}=2.7$. This Bayes factor indicates only anecdotal evidence in favor of the null hypothesis of no effect of mirtazapine over the specified alternative.
@@
Figure 3: Power and sample size calculations for randomized clinical trial assessing the effect of mirtazapine on agitated behavior in patients with dementia (Banerjee et al., , 2021). The Bayes factor assumed for the analysis contrasts $H_{0}\colon\theta=0$ to $H_{1}\colon\theta=-6$ related to the mean difference $\theta$ in CMAI scores between mirtazapine and placebo. The power is computed under either a point prior at the assumed effect under the alternative (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange).
@@
Taking a step back and assuming that no data have yet been collected, we can use equations (3) and (8) to calculate power and sample size. Figure 3 shows power curves based on the Bayes factor providing strong evidence in favor of the alternative ($\mathrm{BF}_{01}<1/10$) in the top plot, and based on the Bayes factor providing strong evidence in favor of the null ($\mathrm{BF}_{01}>10$) in the bottom plot. The colors indicate under which design prior the power was computed.
@@
Focusing on the top plot, we can see from the green curve that at least $n=124$ observations per group are required to ensure a target power of $1-\beta=80\%$ assuming that the same point prior is used in the design as for the analysis (i.e., a point prior at $\mu=-6$). This number increases to $n=195$ when we move to a design prior that incorporates parameter uncertainty (blue curve), i.e., a prior that is still centered around $\mu=-6$ but with a standard deviation of $\tau_{d}=2$. Finally, if we look at the orange power curve computed assuming that the null hypothesis is true ($\theta=0$), we can see that the probability of misleading evidence for the alternative when the null hypothesis is actually true (the type I error rate) is very low and appears to be reasonably controlled by conventional standards (i.e., below 5%) for each of the two samples sizes. This curve could be presented to a trial regulator to demonstrate that the design is adequately calibrated.
@@
Focusing now on the bottom plot, we can see that the sample size $n=124$ based on the point prior also ensures a power of $1-\beta=80\%$ for finding evidence for the null hypothesis. This is due to the symmetric nature of the point versus point hypothesis Bayes factor considered in this example, as swapping the null $\theta_{0}$ and alternative $\mu$ in formula (10) does not change the resulting sample size. Finally, looking at the green and blue curves we can see that the probability of misleading evidence in favor of the null when the data are generated from the non-null design priors seems to be reasonably well controlled (below the conventional 5%) for the point design prior (green), while it is slightly inflated for the normal design prior (blue) for sample sizes larger than 88.
@@ 5.2 Comparison to Monte Carlo simulation methods
Schönbrodt and Wagenmakers, (2018) proposed a Monte Carlo simulation approach for power and sample size calculations for Bayes factor analyses, and provide the R package BFDA (Schönbrodt and Stefan, , 2019) for this purpose. The idea is to simulate data sets under an assumed design prior and sample size, and then analyze each data set with a specified Bayes factor. This results in a distribution of Bayes factors from which the power can be calculated. The simulation is then repeated for other sample sizes until the desired power is achieved. This approach can be used in quite general settings, but can be computationally intensive and comes with Monte Carlo error.
@@
Figure 4: Power and sample size calculation for standardized mean difference example. The Bayes factor assumed for the analysis contrasts $H_{0}\colon\theta=0$ to $H_{1}\colon\theta\neq 0$ with $\theta\mid H_{1}\sim\mathrm{N}(\mu=0,\tau^{2}=1/2)$ prior assigned to the standardized mean difference $\theta$ under the alternative. The power is computed under either a point prior at a medium effect size (green), a normal prior that incorporates additional parameter uncertainty (blue), or under the null hypothesis (orange). For comparison, 1000 Monte Carlo simulations are performed with the BFDA package to obtain a simulation-based power curve approximation (dashed line with one Monte Carlo standard error band).
@@
We will now look at an adaptation of an example examined in Schönbrodt and Wagenmakers, (2018, p. 133). Suppose we want to test the null hypothesis that a standardized mean difference $\theta$ is zero ($H_{0}\colon\theta=0$) versus the alternative hypothesis that it is different from zero ($H_{1}\colon\theta\neq 0$). We assume a $\theta\mid H_{1}\sim\mathrm{N}(0,1/2)$ analysis prior for the standardized mean difference under the alternative, similar to the Cauchy prior with scale $1/\sqrt{2}$ that was assumed by Schönbrodt and Wagenmakers. As they did, we also investigate two design priors: either a point prior at $\mu_{d}=0.5$ , which is a convention for a ‘medium‘ standardized mean difference in psychology (Cohen, , 1992), or a normal prior at $\mu_{d}=0.5$ with standard deviation $\tau_{d}=0.1$ to incorporate parameter uncertainty. Figure 4 shows the resulting power curves computed from equation (6), along with the BFDA Monte Carlo simulation approximations (with Monte Carlo standard error bands) for comparison.
We see that closed-form and simulation curves closely align in all cases. The latter shows a maximum Monte Carlo error of 0.016. This Monte Carlo error may or may not be negligible for practical purposes, as it may lead to slight over- or underestimation of the sample size. For example, in animal studies, an overestimation of the sample size by even one animal may be highly problematic and thus a higher number of simulations may be preferred, whereas in other studies this would not be a problem at all.
@@
Looking at the green curves in the top plot, we see that a sample size of $n=153$ per group is sufficient to achieve a target power of $95\%$ with a Bayes factor threshold of $k=1/6$ under the point design prior. This sample size is slightly larger than the $n=146$ that Schönbrodt and Wagenmakers obtained in their calculations, which instead assumed a Cauchy analysis prior. As in the previous example, incorporating parameter uncertainty via a normal design prior increases the required sample size, in this case to $n=211$. Looking at the orange curve in the bottom plot, we see that these sample sizes only have modest power of around $20\%$ and $50\%$, respectively, for obtaining evidence in favor of the true null hypothesis (at a threshold of $k=6$). To obtain a power of $95\%$, a sample size of $n=6691$ per group would be required (not shown in Figure 4). This illustrates the well-known fact that evidence accumulates slower for the null than the alternative when the Bayes factor involves testing a point null against a composite alternative with normal analysis prior (Johnson and Rossell, , 2010). Finally, looking at the orange curve in the top plot and the green/blue curves in the bottom plot, we see that the probability of misleading evidence in favor of the incorrect hypothesis appears to be adequately controlled, as these curves are virtually constant at zero over the entire range of sample sizes.
Appendix E shows a more systematic comparison with Monte Carlo simulation, including many other design and analysis prior conditions. To verify that the methods work as intended under a given condition, a sample size was first calculated using closed-form (if available) or root-finding methods to ensure a desired target power for a specified Bayes factor threshold. Subsequently, data were simulated based on this sample size and the design prior. Bayes factors were then calculated from the simulated data, and power was estimated empirically from the proportion of Bayes factors below the specified threshold. These simulation-based power estimates were spread closely around the specified target power in all conditions.
@@ 6.1 Informed Bayesian t-test
with $t$ the $t$-test statistic, $n$ the effective sample size (the actual number of observations/pairs for one-sample/paired $t$-tests, or $n=1/(1/n_{1}+1/n_{2})$ for two-sample $t$-tests), $\mathrm{T}_{\nu}(\cdot\mid\mu,\tau)$ the location-scale $t$ density with degrees of freedom $\nu$, location $\mu$, and scale $\tau$, and $\mathrm{NCT}_{\nu}(\cdot\mid\lambda)$ the non-central $t$ density with degrees of freedom $\nu$ and non-centrality parameter $\lambda$ (Johnson et al., , 1995, chapters 28 and 31). The subscript $[a,b]$ denotes truncation of a distribution to the interval from $a$ to $b$, i.e., $f(x)_{[a,b]}=\{f(x)1_{[a,b]}(x)\}/\{F(b)-F(a)\}$. For example, for a one-sided test in positive direction, we have $a=0$ and $b=+\infty$, while for a two-sided test we have $a=-\infty$ and $b=+\infty$. When the prior under the alternative is set to a (scaled) Cauchy distribution ($\kappa=1$, $\mu=0$, $a=-\infty$, $b=+\infty$), the Bayes factor reduces to the ‘Jeffreys-Zellner-Siow’ (JZS) Bayes factor (Jeffreys, , 1939; Zellner and Siow, , 1980), which is often used as a ‘default’ Bayes factor in the social sciences (Rouder et al., , 2009). Setting other values for these parameters allows data analysts to incorporate directionality, such as a one-sided JZS Bayes factor (Wetzels et al., , 2009), or prior knowledge about the standardized mean difference, potentially improving the efficiency of the test (Stefan et al., , 2019). The Bayes factor (13) is also implemented in our package in the function tbf01.
@@
To calculate a sample size that ensures a certain target power, we can again use numerical root-finding. To illustrate this, we now revisit the example from Schönbrodt and Wagenmakers, (2018) which was already re-analyzed in Section 5.2 with a normal analysis prior. We now specify the analysis parameters $\mu=0,\tau=1/\sqrt{2},\kappa=1,a=0,b=+\infty$, i.e., a one-sided JZS Bayes factor with scale $1/\sqrt{2}$ (Wetzels et al., , 2009), as in the original analysis. Using the previously described iterative approach, we calculate that a sample size of $n=143$ is required under a point design prior at $\mu_{d}=0.5$ to achieve a target power of 95%. This differs only slightly from the simulation-based $n=146$ reported by Schönbrodt and Wagenmakers, possibly due to Monte Carlo error.